OpenStudy (anonymous):

find the limit if it exists: lim x->-1 (x^3+1)/(x+1). i think there is no limit for this problem. i can't factor out the numerator

OpenStudy (anonymous):

or am i allowed to just plug in the -1 and solve?

OpenStudy (anonymous):

even so i get a 0 for the denom making it have no limit?

hartnn (hartnn):

there is a limit

hartnn (hartnn):

factor x^3 +1

OpenStudy (cwrw238):

it u plug in x=-1 the expression = 0/0 which is indeterminate

hartnn (hartnn):

\(\huge a^3+b^3=(a+b)(a^2-ab+b^2)\)

OpenStudy (anonymous):

(x+1)(x2-x+1) ?

hartnn (hartnn):

yup, now cancel x+1 from numerator and denominator

OpenStudy (anonymous):

(-1)^2-(-1)+1?

OpenStudy (anonymous):

final answer of 1?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

oh yes, sorry, forgot to change the signs

OpenStudy (anonymous):

no matter

hartnn (hartnn):

so u got 3 ?

OpenStudy (anonymous):

yep

hartnn (hartnn):

\(\large \frak{\color{red}{WELCOME \quad TO \quad OPENSTUDY \quad:D}}\)

OpenStudy (anonymous):

lol

hartnn (hartnn):

this is your first day here, right ? so a warm welcome :)

OpenStudy (anonymous):

:D hey i got a question for another problem. need to double check it. starting a new one!

hartnn (hartnn):

sure , go ahead :)